# IFT-UAM/CSIC-25-21

### v2

@@ Abstract
Building upon the recent findings regarding inverse phase transitions in the early universe, we present the first natural realisation of this phenomenon within a supersymmetry-breaking sector. We demonstrate that inverse hydrodynamics, which is characterized by the fluid being aspirated by the bubble wall rather than being pushed or dragged, is actually not limited to a phase of (re)heating but can also occur within the standard cooling cosmology. Through a numerical analysis of the phase transition, we establish a simple and generic criterion to determine its hydrodynamics based on the generalised pseudo-trace. Our results provide a proof of principle highlighting the need to account for these new fluid solutions when considering cosmological phase transitions and their phenomenological implications.
@@
Introduction – Phase transitions (PTs) in the early universe plasma, usually called cosmological phase transitions, are fascinating phenomena. First order PTs (FOPTs) proceeding via the nucleation and expansion of bubbles of the true vacuum inside a sea of false vacuum are of particular interest as they can be at the origin of the matter-antimatter asymmetry of the universe (baryogenesis) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], lead to the production of dark matter [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] and primordial black holes [28, 29, 30, 31, 32], and can be a powerful source of primordial gravitational waves (GWs) as well [33, 34, 35, 36, 37]. The broad program to discover and investigate a possible background of GWs by current experiments such as Ligo–Virgo–Kagra [38] and Pulsar Timing Arrays [39], as well as future detectors such as the LISA [40] and the Einstein Telescope [41], opens the unique opportunity of probing the existence of FOPTs and of new fundamental physics. Indeed, FOPTs appear naturally in a large variety of scenarios beyond the Standard Model (BSM) like composite Higgs [42, 43, 44, 45, 46], extended Higgs sectors [47, 48, 49, 50, 51, 52, 53, 54], axion models [55, 56], dark Yang-Mills sectors [57, 58], $B-L$ breaking sectors [59, 60] and SUSY breaking sectors [61, 62], and may also be catalyzed by impurities in the early universe, see e.g. [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75], as well as occur in the core of neutron stars [69, 76, 77, 78].
@@
The dynamics of FOPTs involve a non–trivial interplay between the bubble wall and the surrounding plasma, which is pivotal in determining the phenomenology of the PT including the GW emission. The hydrodynamical modes describing the bulk fluid motion in the background of an expanding bubble during a direct FOPT have been classified a long time ago [79, 80, 81, 82]: they consist of detonations and deflagrations, together with a class of hybrid modes interpolating between the two. For all these solutions, the fluid is either dragged or pushed by the bubble wall. The bulk fluid velocity is then always aligned with the wall velocity in the plasma frame [^1] The plasma frame is defined as the frame in which the centre of the bubble is at rest. . In the case of the so–called inverse PTs, the plasma is instead aspirated inside the expanding bubble and the fluid flows in the opposite direction of the bubble wall motion in the plasma frame [83, 84]. These solutions have been so far studied in the context of a (re)heating PT [85, 83, 84], where the temperature of the system increases with time, and thus have been associated to superheated bubbles, see also [86, 87].
In this paper, we show that inverse hydrodynamics is actually not limited to the heating scenario mentioned above, but can instead take place during the standard cooling of the universe, for instance during radiation domination. Interestingly, we find the emergence of this novel hydrodynamics in the context of supersymmetry (SUSY) and $R$–symmetry breaking. In particular, the PT along the pseudomodulus direction within the minimal O’Raifeartaigh model for spontaneous SUSY breaking turns out to be inverse in a relevant part of its parameter space. As we shall see, the occurrence of inverse bubbles can be traced back to the properties of the SUSY spectrum and its distinctive thermal history, opening the possibility of testing these properties through their imprint on the PT and the corresponding GWs. We also provide a precise characterisation of the inverseness of a FOPT by providing a simple criterion based on the sign of a generalised pseudo-trace to readily discriminate between direct and inverse FOPTs. Our findings extend the understanding of inverse phase transitions and their phenomenological implications, offering new avenues for studying their impact on gravitational wave signatures and beyond.
Figure 1: Left panel: Mass spectrum of the SUSY-breaking sector. The fermionic mass eigenstates are shown in red and blue, while the scalar mass eigenstates are depicted in grey and black. The latter are split into pairs around the corresponding fermionic mass eigenstates. Right panel: Tree-level (black) and one-loop corrected (red) potential of the SUSY model along the pseudomodulus direction.
@@
where we have used the standard superspace notation. A thorough characterisation of such PTs within different realisations of the SUSY breaking sector was provided in [62]. In our study, as a minimal benchmark model, we focus on the O’Raifeartaigh model [88]. In addition to the pseudo-modulus, the SUSY breaking sector contains four chiral superfields $\phi_{1},\tilde{\phi}_{1},\phi_{2},\tilde{\phi}_{2}$. The superpotential takes the form
@@
The model in Eq.(2) preserves an $R-$symmetry, which typically accompanies dynamical SUSY breaking [89, 90] and under which $X$ has charge two [^2] One might wonder if the model we are considering is an unrealistic toy model or if it such model could have a phenomenological interest. Indeed, let us emphasize that $R$ symmetry is restored in the zero-temperature phase which is unphysical. However, another sector might be responsible for the low-energy breaking of $R$ symmetry. Actually, such a situation is necessary in the split SUSY [91] and mini-split [92] realisations., $R[X]=2$. The modulus $x$ thus serves as the order parameter governing the spontaneous breaking of the $R$-symmetry. The vacuum expectation values that minimize the tree–level potential are given by $\phi_{1}=\phi_{2}=\tilde{\phi}_{1}=\tilde{\phi}_{2}=0$, leading to a tree-level vacuum energy of $V^{\rm min}_{\rm tree}=|F|^{2}$. Consequently, the potential remains constant as a function of $x$ indicating that supersymmetry is broken irrespective of $x$, while $R$-symmetry is restored at the origin.
On the other hand, the mass spectrum plays a crucial role in shaping the behaviour of the potential for $x$ at the loop level. To analyse this, we diagonalise the scalar and fermion mass matrices, defined as follows
${}\mathcal{M}^{2}_{S,ij}\equiv{1\over 2}\bigg{(}\frac{\partial^{2}V}{\partial\phi_{i}\partial\phi_{j}}\bigg{)}\,,\quad\mathcal{M}_{F,ij}\equiv{1\over 2}\bigg{(}\frac{\partial^{2}W}{\partial\phi_{i}\partial\phi_{j}}\bigg{)}\,.$ (3)
The scalar mass eigenstates are split in pairs around the fermionic mass eigenstates, as we can see in the left panel of Fig.1. There are also light fields from the superfield $X$, corresponding to the pseudomodulus, $x$, the $R$–axion, $a$, and the goldstino, $\tilde{G}$, which will have an influence on the hydrodynamics as they contribute to the number of relativistic species in the plasma. In Appendix A, we present the computation of the effective potential. Due to the one-loop corrections, this acquires a global minimum at the origin, as we can see in the right panel of Fig.1. A remarkable property of this potential, which is a reflection of the underlying SUSY, is that even at one-loop level it remains extremely flat away from the origin.
@@
However, finite temperature effects (described in Appendix A) break SUSY explicitly and can have a strong impact on the pseudomodulus effective potential. The typical thermal history of the minimal O’Raifeartaigh model considered here is then as follows [62, 93, 94]: at high temperatures, $T\gtrsim\sqrt{F}$, the system has a single vacuum state with $\langle x\rangle=0$. At lower temperatures, a new local minimum of the effective potential appears at relatively large field values, $\langle x\rangle/\sqrt{F}\gg 1$, which can become the true vacuum of the theory below a certain critical temperature, $T_{c}$. This state with broken $R$-symmetry will however become metastable and then eventually disappear at even lower temperatures, given that the only minimum at $T=0$ is at $\langle x\rangle=0$, as already shown in the right panel of Fig. 1.
Figure 2: The free energy difference between the $R$–symmetry breaking and preserving vacua as a function of the temperature. For $T>T_{c}$, the symmetric vacuum is the global one, while for $T<T_{c}$, the broken phase with $\langle x\rangle\neq 0$ becomes favoured. The nucleation temperature (dashed red line) is found to be close to the temperature where the barrier disappears, $T/\sqrt{F}\approx 0.66$. At lower temperatures, $T/\sqrt{F}\lesssim 0.6$, the $R$–symmetry preserving vacuum becomes again the global minimum. The condition $D\vartheta=0$ marks the change from direct to inverse hydrodynamics. The insets indicate a sketch of the free energy, including the position of the relevant vacua in the corresponding temperature range.
In Fig. 2 we present an example of such thermal history for a characteristic benchmark point with $m/\sqrt{F}=2$ and $\lambda=1.67$, by showing the free energy difference between the vacuum at $\langle x\rangle=0$ and $\langle x\rangle\neq 0$, for temperatures where they both exist and are classically stable. At $T=T_{c}$ the free energy difference changes sign indicating that the minimum with broken $R$–symmetry starts to be favoured, and the system can now tunnel from the $R$–symmetry preserving vacuum to this new phase. This PT is first order and is controlled by a thermal barrier that will eventually disappear at $T/\sqrt{F}\simeq 0.66$. Nucleation in the expanding universe will in fact take place at a slightly higher temperature, $T_{\text{nuc}}/\sqrt{F}\approx 0.665$, as evaluated numerically and shown by the dashed red line. As the system continues to cool, the barrier reappears, and the symmetry-breaking minimum is gradually lifted until it once again becomes degenerate with the origin at $T=T_{c,2}$. Finally, at very low temperatures, the symmetry breaking minimum disappears and the symmetry is restored again, causing a second FOPT.
@@
In this paper, we will focus on the first transition that will take place in the expanding universe, namely the $R$–symmetry breaking FOPT towards the vacuum with $\langle x\rangle\neq 0$. As we shall see, this PT can actually proceed according to either the direct or the inverse hydrodynamics (see Ref. [83, 84]) depending on the microscopic coupling constant $\lambda$ entering the superpotential in Eq. (2) (while the second FOPT restoring $R$–symmetry will always be direct). This is shown in Fig. 3, where the nucleation temperature as a function of $\lambda$ is indicated by the red line (keeping $m/\sqrt{F}=2$ fixed). As we can see, for all points nucleation takes place very close to the temperature where the barrier actually disappears. For $\lambda\lesssim 1.63$, bubble nucleation occurs in the region where the hydrodynamics will be the one based on the direct detonation and deflagration type of solutions, while for $1.63\lesssim\lambda\lesssim 1.68$ the hydrodynamics will be inverse, as we shall discuss in detail in the next sections.
For each value of the coupling constant $\lambda$, we also evaluate numerically the duration of the FOPT compared to one Hubble time, obtaining typical values of $\beta/H=\mathcal{O}(10^{4})$, where as usual $\beta=-\frac{d}{dt}S_{3}/T$ and $S_{3}$ is the $O(3)$ bounce action (see Appendix B for more details).
@@
Thermodynamics and hydrodynamics of the SUSY model – In the early universe, FOPTs can be modelled as the interplay between a scalar field $\phi$, whose vacuum expectation value represents the order parameter of the transition, and the surrounding plasma which is often well described by a relativistic fluid. The energy–momentum tensor of the system consists then of those two contributions, $T^{\mu\nu}=T^{\mu\nu}_{\text{fluid}}+T^{\mu\nu}_{\phi}$, with
$T^{\mu\nu}_{\phi}$ $=\partial^{\mu}\phi\partial^{\nu}\phi-g^{\mu\nu}\left(\frac{1}{2}(\partial\phi)^{2}-V(\phi)\right)$ (4a)
@@
The conservation of the energy–momentum tensor across the phase boundary, $\nabla_{\mu}T^{\mu\nu}=0$, gives the following junction conditions for the quantities that are discontinuous at the bubble wall
$w_{+}\gamma_{+}^{2}v_{+}$ $=w_{-}\gamma_{-}^{2}v_{-}\ ,$ (6a)
$w_{+}\gamma_{+}^{2}v_{+}^{2}+p_{+}$ $=w_{-}\gamma_{-}^{2}v_{-}^{2}+p_{-}\ ,$ (6b)
@@
where the subscript “$\pm$” denotes quantities in front of/behind the phase boundary, so that for instance “--” always represents the interior of the bubble. To be explicit, $w_{+}=w_{s}(T_{+})$, $w_{-}=w_{b}(T_{-})$ (and similarly for $p_{\pm}$), where the label “$s/b$” denotes the symmetric/broken phase of the $R$–symmetry in the case of interest.
Upon rearranging the junction conditions, we arrive at the familiar relations for the fluid velocities ahead and behind the wall, the energies and the pressures [^3] We remind that the velocities $v_{\pm}$ have to be understood in the front frame, where the bubble wall is at rest.
${}v_{+}v_{-}=\frac{p_{+}-p_{-}}{e_{+}-e_{-}}\ ,\qquad\frac{v_{+}}{v_{-}}=\frac{e_{-}+p_{+}}{e_{+}+p_{-}}\ .$ (7)
Figure 3: The nucleation temperature (red) is obtained as a function of $\lambda$ by numerically solving the condition $S_{3}/T=140$, which corresponds to setting $\sqrt{F}\sim$ TeV for concreteness. The blue-shaded region indicates the region of the inverse FOPTs, which occur approximately within the window $\lambda\in[1.63,1.675]$, while the white region indicates the parameter space where the FOPT will be direct. We also show the boundary between inverse and direct FOPTs according to the criteria $D\vartheta=0$ and $D\theta=0$, see discussion around Eq. (8), which agree well with the explicit evaluation of the hydrodynamics. For this figure we fixed $m/\sqrt{F}=2$.
Let us now examine the possible hydrodynamics of the $R$–symmetry breaking PT. The junction conditions above can be solved numerically by referring to the pressure and energy densities as evaluated directly from the free energy within our particle physics model. The allowed values for the $(v_{-},v_{+})$ pairs are shown in Fig. 4 for a representative benchmark point. The matching conditions in Eq. (7) are solved for $v_{\pm}$ in terms of the temperatures ahead and behind the wall, $T_{\pm}$. For consistency, we restrict $T_{+}$ to lie between $T_{c}$ and the temperature when the barrier disappears, $T_{\rm no\ barrier}$, as this is the range for which the FOPT can actually take place. The various $v_{\pm}$ trajectories in Fig. 4 are then shown together with the corresponding temperature $T_{+}$ according to the colour code. Because of the consistency condition on $T_{+}$ and the properties of our system free energy, the branches do not populate the entire $v_{\pm}\in(0,1)$ parameter space. The regions corresponding to inverse and direct hydrodynamics are indicated by solid and dashed lines, respectively. We find that these regions remain neatly separated across the entire $(v_{-},v_{+})$ plane, except for a small overlap in the regime of hybrid solutions (bottom-right corner). More details are presented in Appendix C. As a comparison, a similar discussion of the inverse branches in the case of the (template) $\mu\nu$-model is provided in Appendix D.
If we further specify the temperature of the FOPT as the one evaluated numerically from the bubble nucleation condition  [^4] The relation $T_{\text{nuc}}=T_{+}$ only holds for detonations and anti–deflagrations. For the other expansion modes, we still use this as a sensible approximation to identify the relevant $v_{\pm}$ branch., we can select the bright red branch as the relevant one for this specific benchmark point  [^5] Notice that, as the matching conditions can not uniquely determine the bubble wall velocity, the actual value of $v_{\pm}$ cannot be determined by hydrodynamics only and the full red branch can in principle be realised.. As we can see, the FOPT occurs in the inverse hydrodynamic regime, as anticipated in Fig. 3.
Figure 4: Possible solutions to the fluid matching conditions for $(v_{-},v_{+})$ for the $R$–symmetry breaking FOPT under consideration, plotting the relevant branches for different values of $T_{+}\in[T_{\rm no\ barrier},T_{c}]$. Dashed lines correspond to direct phase transitions, while solid lines indicate inverse transitions, determined by the sign of $\alpha_{\vartheta}$. The solid red line highlights the relevant branch at the nucleation temperature. The red-shaded area marks the region of strong (inverse) detonations and strong (inverse) deflagrations. In the bottom right corner, a zoomed-in view of the hybrid solution region reveals an overlap between different branches (see Appendix C for more details).
@@
We defined inverse PTs as transitions displaying negative bulk velocities in the plasma frame: rather than being pushed outward, the surrounding plasma is drawn inward, effectively being aspirated into the expanding bubble. Let us now provide a sharper characterisation, or criterion, of inverse hydrodynamics which extends the intuitive one put forward in Ref. [83], according to which inverse PTs are found when the transition proceeds against the vacuum energy (namely, the $T=0$ effective potential for the order parameter). We find that a fully general characterization of inverse hydrodynamics can be obtained by defining a generalised pseudo-trace, $\alpha_{\vartheta}$, which extends the definition within the bag EoS adopted in  [83] as well as the pseudo-trace, $\alpha_{\theta}$, introduced in  [95]
@@
where the $D$ and $\delta$ are defined as $Df=f_{+}(T_{+})-f_{-}(T_{+})$ and $\delta f=f_{-}(T_{+})-f_{-}(T_{-})$. For given values of $T_{\pm}$, which as discussed below (7) can be related to $v_{\pm}$ via the matching conditions, inverse hydrodynamics takes place for $\alpha_{\vartheta}<0$, while the standard one is realised for $\alpha_{\vartheta}>0$. This criterion is the one used to identify inverse and direct branches in Fig. 4. In this way, we discover that PTs proceeding against the vacuum energy can nonetheless display direct hydrodynamics.
@@
Notice that for relatively weak PTs with $T_{+}\simeq T_{-}$, $\delta e/\delta p\simeq 1/c_{s,-}^{2}$, with $c_{s,-}$ being the speed of sound in the broken phase, and Eq. (8) reduces to $\alpha_{\theta}$ as defined in Ref. [95]. In the special case of a strictly constant speed of sound, this definition further reduces to $\alpha_{\theta}$ as derived in the template $\mu\nu$ model, see Eq. (52) in Appendix D. Finally, when the speed of sound is $c_{s}^{2}=1/3$ as for a relativistic gas, this definition reduces to $\alpha_{+}$ as considered in Ref. [83].
@@
One can show that FOPTs with $\alpha_{\vartheta}=0$ represent the limit of weak hydrodynamics, where $\Delta e=0$ and $\Delta p=0$, with $\Delta f=f_{+}(T_{+})-f_{-}(T_{-})$. By continuity, this is supposed to separate inverse from direct PTs. This is confirmed by the results shown in Fig. 3, where the line of vanishing $\alpha_{\vartheta}$ actually corresponds to the boundary between direct and inverse regions, which are determined independently by solving the fluid equations. As we can see, the approximate condition in terms of the pseudo-trace, $\alpha_{\theta}=0$, reproduces this separation fairly well. This can be traced back to the fact that the speed of sound is not strongly temperature dependent in this model. We present the precise relation between the generalised pseudo-trace and the pseudo-trace in the $\mu\nu$ model in Appendix D.

### v3

@@ Abstract
Building upon the recent findings regarding inverse phase transitions in the early universe, we present the first natural realisation of this phenomenon within a supersymmetry-breaking sector. We demonstrate that inverse hydrodynamics, which is characterized by the fluid being sucked by the bubble wall rather than being pushed or dragged, is actually not limited to a phase of (re)heating but can also occur within the standard cooling cosmology. Through a numerical analysis of the phase transition, we establish a simple and generic criterion to determine its hydrodynamics based on the generalised pseudo-trace. Our results provide a proof of principle highlighting the need to account for these new fluid solutions when considering cosmological phase transitions and their phenomenological implications.
@@
Introduction – Phase transitions (PTs) in the early universe plasma, usually called cosmological phase transitions, are fascinating phenomena. First order PTs (FOPTs) proceeding via the nucleation and expansion of bubbles of the true vacuum inside a sea of false vacuum are of particular interest as they can be at the origin of the matter-antimatter asymmetry of the universe (baryogenesis) [85, 106, 98, 40, 43, 92, 34, 33, 35, 96, 13, 70, 17, 42], lead to the production of dark matter [57, 20, 68, 14, 19, 9, 93, 21, 11, 18, 79, 63, 12] and primordial black holes [80, 78, 75, 64, 89], and can be a powerful source of primordial gravitational waves (GWs) as well [110, 67, 82, 81, 76]. The broad program to discover and investigate a possible background of GWs by current experiments such as LIGO–Virgo–Kagra [104] and Pulsar Timing Arrays [32], as well as future detectors such as the LISA [39] and the Einstein Telescope [94], opens the unique opportunity of probing the existence of FOPTs and of new fundamental physics. Indeed, FOPTs appear naturally in a large variety of scenarios beyond the Standard Model (BSM) like composite Higgs [101, 15, 58, 103, 59], extended Higgs sectors [51, 84, 109, 10, 60, 8, 16, 25], axion models [52, 108], dark Yang-Mills sectors [66, 95], BLB-L breaking sectors [74, 1] and SUSY breaking sectors [47, 48], and may also be catalysed by impurities in the early universe, see e.g. [107, 69, 111, 97, 37, 72, 23, 3, 30, 2, 73, 31, 29], as well as occur in the late Universe in the core of neutron stars [23, 22, 41, 24].
@@
The dynamics of FOPTs involve a non–trivial interplay between the bubble wall and the surrounding plasma, which is pivotal in determining the phenomenology of the PT including the GW emission. The hydrodynamical modes describing the bulk fluid motion in the background of an expanding bubble during a direct FOPT have been classified a long time ago [87, 83, 86, 56]: they consist of detonations, hybrids and deflagrations. For all these solutions, the fluid is either dragged or pushed (or both) by the bubble wall. The bulk fluid velocity is then always aligned with the wall velocity in the plasma frame [^1] The plasma frame is defined as the frame in which the centre of the bubble is at rest. . In the case of the inverse PTs, the plasma is instead sucked inside the expanding bubble and the fluid flows in the opposite direction of the bubble wall motion [26, 27] (see also Ref. [50] for the case of droplet collapse). These solutions have been so far studied in the context of a (re)heating PT [36, 26, 27] where the temperature of the system increases with time, and thus have been associated to superheated bubbles, see also [38, 53].
In this Letter, we show that inverse hydrodynamics is actually not limited to the heating scenario mentioned above, but can instead take place during the standard cooling of the universe as well. Remarkably, we find the emergence of this novel hydrodynamics during a seemingly standard, symmetry–breaking phase transition. This takes place within the context of dynamical supersymmetry (SUSY) and RR–symmetry breaking, which represents the first explicit example for this class of FOPTs. Our findings extend the current understanding of what types of cosmological phase transitions can actually take place, thus opening up new directions for studying the corresponding GW signatures and other phenomenological aspects.
@@
where we have used the standard superspace notation. In our study, as a minimal benchmark model, we focus on the O’Raifeartaigh model [100]. In addition to the pseudo-modulus, the SUSY breaking sector contains four chiral superfields $\phi_{1},\tilde{\phi}_{1},\phi_{2},\tilde{\phi}_{2}$. The superpotential takes the form
@@
The model preserves a global $U(1)$ $R$-symmetry, which typically accompanies dynamical SUSY breaking [99, 71], and under which $X$ has charge two, $R[X]=2$. The vacuum expectation value (vev) of $x$ is the order parameter for spontaneous $R$-symmetry breaking, while the additional scalar fields from $\phi_{i}$ and $\tilde{\phi}_{i}$ will have vanishing vev in all phases. The tree–level vacuum energy is $V^{\rm min}_{\rm tree}=|F|^{2}$ with $x$ being a flat direction, indicating that supersymmetry is broken irrespectively of $\langle x\rangle$ while $R$-symmetry is preserved only at the origin, $\langle x\rangle=0$.
The pseudomodulus flat direction is however lifted at the loop level, and the shape of the potential for $x$ is controlled by the mass spectrum of the theory. One can see that this includes massive particles from the $\phi_{i}$ and $\tilde{\phi}_{i}$ superfields, with the scalar eigenstates split in pairs around the fermion ones, as well as massless fields from the superfield $X$ corresponding to the pseudomodulus, $x$, the RR–axion, $a$, and the goldstino, $\tilde{G}$. At one–loop, the potential for $x$ acquires a global minimum at the origin, while remaining remarkably flat at large field values as a reflection of the underlying SUSY.
@@
Finite–temperature effects, on the other hand, break SUSY explicitly and have a strong impact on the pseudomodulus effective potential. The typical thermal history of the minimal O’Raifeartaigh model considered here is then as follows [48, 46, 77]: at very high temperatures, $T\gtrsim\sqrt{F}$, the system has a single vacuum state, $\langle x\rangle=0$, and RR–symmetry is preserved. At lower temperatures, a new local minimum of the effective potential appears at relatively large field values, $\langle x\rangle/\sqrt{F}\gg 1$, which becomes the true vacuum of the theory below a certain critical temperature, $T_{c}$. This vacuum with broken $R$-symmetry will however become metastable and eventually disappear at even lower temperatures, given that the only minimum at zero temperature is at $\langle x\rangle=0$.
Overall, the system undergoes two phase transitions, namely (1) the breaking of the RR–symmetry at high temperatures and (2) its restoration at low temperatures, which turn out to be first order and governed by a thermal barrier. More details on the standard derivation of the effective potential for $x$ and the associated thermal history can be found in App. C and references therein [45, 102, 49], as well as in Ref. [48].
@@
In this paper, we will focus on the first transition that will take place in the expanding universe, namely the RR–symmetry breaking FOPT: $\langle x=0\rangle\rightarrow\langle x\neq 0\rangle$. As it turns out, this FOPT can actually proceed according to either the direct or the inverse hydrodynamics (the latter presented in Ref. [26, 27]) depending on the microscopic coupling constant $\lambda$ entering the superpotential in Eq. (2), while the second RR–symmetry restoring FOPT will always be direct.
@@
Thermodynamics and hydrodynamics of RR–symmetry breaking – In the early universe, FOPTs can be modelled as the interplay between a scalar field $\phi$, whose vacuum expectation value represents the order parameter of the transition, and the surrounding plasma which is often well described by a relativistic fluid. The energy–momentum tensor of the system consists then of those two contributions, $T^{\mu\nu}=T^{\mu\nu}_{\text{fluid}}+T^{\mu\nu}_{\phi}$, with
$T^{\mu\nu}_{\phi}$ $=\partial^{\mu}\phi\partial^{\nu}\phi-g^{\mu\nu}\left(\frac{1}{2}(\partial\phi)^{2}-V(\phi)\right)\,$, (3a)
@@
The conservation of the energy–momentum tensor across the phase boundary, $\nabla_{\mu}T^{\mu\nu}=0$, gives the following relations between the velocities, the energies and the pressures  [^2] We remind that the velocities $v_{\pm}$ have to be understood in the front frame, where the bubble wall is at rest.
${}v_{+}v_{-}=\frac{p_{+}-p_{-}}{e_{+}-e_{-}}\ ,\qquad\frac{v_{+}}{v_{-}}=\frac{e_{-}+p_{+}}{e_{+}+p_{-}}\,$, (5)
@@
where the subscript $\pm$” denotes quantities in front of/behind the phase boundary, so that for instance “-” always represents the interior of the bubble.
@@
We defined inverse PTs as transitions displaying negative bulk velocities in the plasma frame: rather than being pushed outward, the surrounding plasma is drawn inward, effectively being sucked into the expanding bubble. Let us now provide a sharper characterization, or criterion, of inverse hydrodynamics which extends the intuitive one put forward in Ref. [26], according to which inverse PTs are found when the transition proceeds against the vacuum energy (namely, the $T=0$ effective potential for the order parameter). We find that a fully general characterization of inverse hydrodynamics can be obtained by defining a generalised pseudo-trace, $\alpha_{\vartheta}$, which indicates the strength of the phase transition and extends the definition within the bag EoS adopted in [26] as well as the pseudo-trace, $\alpha_{\theta}$, introduced in [62],
@@
where the $D$ and $\delta$ are defined as $Df=f_{+}(T_{+})-f_{-}(T_{+})$ and $\delta f=f_{-}(T_{+})-f_{-}(T_{-})$. For given values of $T_{\pm}$, they can be related to $v_{\pm}$ via the matching conditions in (5), then, inverse hydrodynamics takes place for $\alpha_{\vartheta}<0$, while the standard one is realised for $\alpha_{\vartheta}>0$. In this way, we discover that PTs proceeding against the vacuum energy can nonetheless display direct hydrodynamics.
@@
Notice that for relatively weak PTs with $T_{+}\simeq T_{-}$, $\delta e/\delta p\simeq 1/c_{s,-}^{2}$, with $c_{s,-}$ being the speed of sound in the broken phase, Eq. (6) reduces to $\alpha_{\theta}$ as defined in Ref. [62]. In the special case of a strictly constant speed of sound, one can refer to the template $\mu\nu$ model as introduced in Ref. [88] to capture deviations from the relativistic fluid with $c_{s}^{2}\neq 1/3$. In this case, our definition further reduces to $\alpha_{\theta}$ as derived within this template. Finally, when the speed of sound is $c_{s}^{2}=1/3$ as for a relativistic gas, this definition reduces to $\alpha_{+}$ as considered in Ref. [26].
One can show that FOPTs with $\alpha_{\vartheta}=0$ represent the limit of weak hydrodynamics, where $\Delta e=0$ and $\Delta p=0$, with $\Delta f=f_{+}(T_{+})-f_{-}(T_{-})$. By continuity, this is supposed to separate inverse from direct FOPTs.
Figure 1: Possible solutions to the fluid matching conditions for $(v_{-},v_{+})$ for the RR–symmetry breaking FOPT under consideration, plotting the relevant branches for different values of $T_{+}$ between $T_{c}$ and the temperature where the barrier disappears. Dashed lines correspond to direct phase transitions, while solid lines indicate inverse transitions, as determined by the sign of $\alpha_{\vartheta}$. The solid red line highlights the relevant branch at $T_{\text{nuc}}$. The red-shaded area marks the region of strong (inverse) detonations and strong (inverse) deflagrations. In the bottom right corner, a zoomed-in view of the hybrid solution region reveals an overlap between different branches (see Appendix B for more details).
Let us now examine the possible hydrodynamics of the RR–symmetry breaking FOPT. The junction conditions above can be solved numerically by referring to the pressure and energy densities as evaluated directly from the free energy within our particle physics model. The allowed values for the $(v_{-},v_{+})$ pairs are shown in Fig. 1 for a representative benchmark point. The matching conditions in Eq. (5) are solved for $v_{\pm}$ in terms of the temperatures ahead and behind the wall, $T_{\pm}$. For consistency, we restrict $T_{+}$ to lie between $T_{c}$ and the temperature when the barrier disappears, as this is the range for which the FOPT can actually take place. The various $v_{\pm}$ trajectories in Fig. 1 are then shown together with the corresponding temperature $T_{+}$ according to the colour code. Because of the consistency condition on $T_{+}$ and the properties of our system free energy, the branches do not populate the entire $v_{\pm}\in(0,1)$ parameter space. The regions corresponding to inverse and direct hydrodynamics, according to the sign of $\alpha_{\vartheta}$, are indicated by solid and dashed lines, respectively. We find that these regions remain neatly separated across the entire $(v_{-},v_{+})$ plane, except for a small overlap in the regime of hybrid solutions (bottom-right corner). As a comparison, a similar discussion of the inverse branches in the case of the simplified (template) $\mu\nu$-model is provided in Appendix A, where we find qualitative agreement with the full numerical study of the SUSY model presented here.
In the early universe, bubbles are efficiently formed when the nucleation rate catches up with the Hubble expansion. This condition, presented in more detail in App. D and Refs. [44, 90, 91, 54, 65, 55], connects the onset of the FOPT with a certain nucleation temperature, $T_{\rm{nuc}}$. If we then further specify the temperature of the FOPT as $T_{\rm{nuc}}$ [^3] The relation $T_{\text{nuc}}=T_{+}$ only holds for detonations and anti–deflagrations. For the other expansion modes, we still use this as a sensible approximation to identify the relevant $v_{\pm}$ branch., we can select the bright red branch as the relevant one for this specific benchmark point. Notice that, as the matching conditions can not uniquely determine the bubble wall velocity, the actual value of $v_{\pm}$ cannot be pinned down by the hydrodynamics only, and the full red branch can in principle be realised. On the other hand, when taking the wall velocity as an additional input, the fluid profile can be fully determined. As we can see, the FOPT within this benchmark point occurs in the inverse hydrodynamic regime.
In Fig. 2 we perform a scan over the model parameter space, by fixing $m/\sqrt{F}=2$ and varying the coupling constant $\lambda$. The red line indicates the nucleation temperature, which always happens to be very close to the temperature where the barrier actually disappears.
@@
For $\lambda\lesssim 1.63$, bubble nucleation occurs in the region where the hydrodynamics will be the one based on the (direct) detonation and deflagration types of solutions, while for $1.63\lesssim\lambda\lesssim 1.68$ the hydrodynamics will be inverse. We can also notice that the condition of vanishing $\alpha_{\vartheta}$ actually corresponds to the boundary between direct and inverse regions, which are determined independently by solving the fluid equations. As we can see, the approximate condition in terms of the pseudo-trace, $\alpha_{\theta}=0$, reproduces this separation fairly well. This can be traced back to the fact that the speed of sound is not strongly temperature dependent in this model.
Figure 2: The nucleation temperature (red line) is obtained as a function of $\lambda$ by numerically solving the condition $S_{3}/T=140$, which corresponds to setting $\sqrt{F}\sim$ TeV for concreteness (see App. D). The blue-shaded (white) region indicates the occurrence of the inverse (direct) FOPTs, whose boundary is shown according to the criteria $D\vartheta=0$ and $D\theta=0$. For this figure we fixed $m/\sqrt{F}=2$.
Figure 3: Free energy of the system at finite temperature evaluated at one–loop at the nucleation temperature, with the arrow indicating the direction of the phase transition towards the minimum with a non–zero $x$ (left panel), together with a sketch of the expanding bubble and its velocity in red, and the fluid profile in green (central panel) for a characteristic inverse detonation. The actual fluid profile in the plasma frame is shown in the right panel. Due to its inverse nature, the fluid velocity is always negative. See main text and Appendix B for details.